OpenStudy (anonymous):

What is the y-value of sin(x) when x= -360°?

OpenStudy (anonymous):

For what x-value(s) does cos(x) = 0?

OpenStudy (anonymous):

@JFraser

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (solomonzelman):

so for very first question you are dealing with a function \(\large\color{black}{{\rm y}=\sin(x)}\) when \(x=360\) what do you get for \( \rm y\) ? Plug in 360 instead of \(x\) into the function.

OpenStudy (solomonzelman):

do you know what \(\sin(360)\) is equivalent to?

OpenStudy (anonymous):

0 @SolomonZelman

OpenStudy (anonymous):

wait 360 is negative.. @SolomonZelman

OpenStudy (solomonzelman):

sin(360)=0 you were correct

OpenStudy (solomonzelman):

now, for the second question you need to find all solutions cos(x)=0 there is an infinity of these solutions, but they all follow certain pattern.

OpenStudy (anonymous):

@Butterfield1215

OpenStudy (solomonzelman):

I will give an example of how to generate this pattern. --------------------------------------------- ((here goes the question)) `What values of x are solutions sin(x)=0 ?` ((here is the solutions)) At first, we know that \(\sin(0)=0\), and therefore \(x=0\) is one possible solution. then there are all co-terminal angles (each \(360^\circ\) you will come to the same value, because the circle/cycle restarts).  so, every +360 degrees (anticlockwise) or every -360 (going clockwise) you meet a condition of sin(x)=0. So, your pattern is \(\large\color{black}{ \displaystyle x= k\times 360;~~~k\in{\bf Z} }\)